OpenStudy (anonymous):

Factor completely 121x^2 - 44x + 4 Choose one answer. a. (11x - 2)2 b. (2x - 11)2 c. (11x - 2)(11x + 2) d. None of the above

OpenStudy (anonymous):

D again?

OpenStudy (anonymous):

121x^2 - 44x + 4 = 0 USE FORMULA ( a^2 - 2ab + b^2) = ( a - b) ^2 Here a = 11x b = 2 ( 11x - 2) ^2 = 0 So 11x = 2 x = 2/11 ANSWER

OpenStudy (anonymous):

So it be A then?

OpenStudy (anonymous):

Yes A seems about right

OpenStudy (anonymous):

Ahh thank you guys! I love algebra but I hate these factoring ones..

OpenStudy (anonymous):

( 11x - 2) ^2